Circular No. 7720

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001el IN NGC 1448 Liebert A. G. Monard, situated 40 km east of Pretoria, South Africa,

reports his visual discovery of an apparent supernova (mag 14.5) on Sept. 17.064 UT. Using a 0.32-m reflector, Monard estimated the new object to be ∼ 9′′ west and 15′′ north of the center of NGC 1448, and he observed SN 2001el again at m_{v} = 14.3 on Sept. 17.930; nothing was noticed at this location to mag 14.6 on Aug. 25.96. G. Bock, Edens Landing, Queensland, estimates mag ∼ 13.2 for SN 2001el from an unfiltered CCD image taken on Sept. 18.61. A. Wassilieff, Palmerston North, New Zealand, reports the following position for SN 2001el from an unfiltered CCD image taken through clouds on Sept. 18.636 (when the new object appeared at mag 13.9) with a 0.38-m reflector: α = 3^{h}44^{m}30^{s}.60, δ = −44^{o}38′23′′.4 (equinox 2000.0), which is 14′′ west and 20′′ north of the nucleus of NGC 1448. J. Biggs provides the following position end figures for SN 2001el from three unfiltered CCD images obtained on Sept. 19.62 with the 0.25-m Mike Candy Telescope at Perth Observatory: 30^{s}.57 ± 0^{s}.02, 23′′.7 ± 0′′.3 (USNO-A2.0 reference stars), which is ∼ 22′′ west and 19′′ north of the center of NGC 1448; the magnitude of the new star was R = 14.5 ± 0.5. SN 1983S also appeared in NGC 1448 (cf. IAUC 3877).

COMET P/2001 R1 (LONEOS) The following improved orbital elements (from additional astrometry

published on MPEC 2001-S05, including prediscovery observations on Aug. 19 by LINEAR) by B. G. Marsden indicate that this comet will pass only 0.014 AU from Mars on 2002 Jan. 10.7 TT, as first suggested by C.-I. Lagerkvist (Uppsala) and G. Hahn (German Aerospace Center, Berlin).

T = 2002 Feb. 17.5491 TT  ω = 24.^{o}7215
e = 0.608441  Ω = 35.5016 }2000.0
q = 1.360575 AU  i = 7.0396
a = 3.474767 AU n^{o} = 0.1521653 P = 6.477 years

COMET P/2001 Q2 (PETRIEW) Total-visual-magnitude estimates: Aug. 22.11 UT, 9.5 (M. V. Zanotta,

Colle Nivolet, Parco Nazionale Gran Paradiso, Italy, 25×150 binoculars); 28.09, 9.2 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); Sept. 5.79, 10.2 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector).

2001 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green